OpenStudy (anonymous):

can someone help me evaluate the expression and write the result in the form a + bi. 7 + 16i/ 8i

OpenStudy (whpalmer4):

Is that \[\frac{7+16i}{8i}\]? If so, a useful tactic is to multiply numerator and denominator by the conjugate of the denominator. If the denominator is \((a+bi)\), the conjugate will be \((a-bi)\). After the dust settles, don't forget that \(i^2=-1\).

OpenStudy (anonymous):

@whpalmer4 yes!

OpenStudy (anonymous):

@whpalmer4 so how exactly do you solve this

OpenStudy (whpalmer4):

I told you... \[\frac{7+16i}{8i}*\frac{-8i}{-8i} = \]

OpenStudy (whpalmer4):

\[\frac{7+16i}{8i}*\frac{-8i}{-8i} = \frac{-8i(7+16i)}{-64i^2} = \frac{-8i(7+16i)}{64} \] Bring it home!